OpenStudy (anonymous):

What are the solutions of the equation 2x2 = 2?

OpenStudy (anonymous):

I'm afraid the question seems somewhat invalid.

OpenStudy (anonymous):

In a a way yes, hold on.

OpenStudy (anonymous):

Just the solutions of 2x^2 =2 would be fine.

OpenStudy (anonymous):

On simplification, isn't it x^2=1 that is, x^2-1=0?

OpenStudy (anonymous):

So, there are no solutions?

OpenStudy (anonymous):

You asked a similar question... x^2-x+x-1=0 that is x(x-1)+(x-1)=0 that is (x+1)(x-1)=0

OpenStudy (anonymous):

2x * 2 = 2 -- -- 2 2 2x =0 - - 2 2 x =0

OpenStudy (anonymous):

not 100% but pretty sure hehe

OpenStudy (anonymous):

Michaelc96. I solved it. ^

OpenStudy (anonymous):

that's wrong nevermind

OpenStudy (anonymous):

My solution, you mean to say?

OpenStudy (anonymous):

no my solution is wrong 2x * 2 = 2 -- -- 2 2 2x =1 - - 2 2 x =.5 but I think that's still wrong

OpenStudy (anonymous):

Yes. Quite. Care to look over and comprehend my solution?

OpenStudy (anonymous):

*I shall go to sleep in the meantime, by the way*

OpenStudy (anonymous):

umm I don't understand it when you come back tell why my work is wrong

OpenStudy (anonymous):

Alright. So what we have here, is 2x^2=2 I divide both sides of my equation by 2. That leaves me x^2=1 Now, I subtract 1 from both sides of my equation. That gives me, x^2-1=1-1 Which is x^2-1=0 Now in both the equations I add +x-x. The expression is 0, which is how I represent in the Right side of the eqution, but on the left side, it is seen as x^2+x-x-1=0 that is, x(x+1)-(x+1)=0 that is, x(x+1)+(-1)(x+1)=0 Now, on the left side of the equation, I take x+1 common for both x, and -1 That gives me, (x-1)(x+1)=0 This is the equation factorised. Post this, I want a solution. Since I have left side is equal to zero, either of the two expressions, x-1=0 or x+1=0. First, I take x-1=0 I add 1 to both sides, leaves me x=1 Now, for x+1=0 I subtract 1 from both sides, that gives me x=-1 So, my solution is that x is either 1 or -1. I hope it is clear. =)